

********************************************************************
UARS WEEKLY STATUS REPORT - 03 APRIL 1998
********************************************************************
Note from Skip:
I'd like to thank everyone associated with UARS for all the thoughts,
calls, cards, flowers, and the HUGE card signed by the Science Team at
the meeting in Pasadena. I returned to work yesterday part time and
I'm getting better every day. The doctors say I'll be back to normal
shortly (as soon as they can figure out what "normal" is in my case.)
...........Skip Reber
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data this week without apparent problems.
***********
PEM
***********
PEM is currently off.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on Mar 26, in the reduced 85 watt power
mode. MLS instrument operations have continued flawlessly. MLS was powered
off at 089:01:38 UT. MLS is scheduled to begin north viewing operations
starting Apr 15.
***********
HRDI
***********
HRDI is in low-power standby mode.
***********
WINDII
***********
WINDII is awaiting activation on April 4, running through April 7.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
All level 0 data for power outage days were received by 27-March-1998.
All current day production processing is up to date.
The CDHF has received a total of 381 CD-ROM requests as of 3-April-1998.
There was one order received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1074C=305L)
84 days MLS (4.90) PROD (1297R-1223C-17F=57L)
5 days CLAES (7.0) ENG (38R-37C=1L)
1 days HAL (4.80) ENG (5R-5C=0L)
The CDHF attended a meeting with various contractors, tele-com to MSFC
and AMES, regarding NISN support of the UARS network. The CDHF provided
attendees original and current network requirements, copies of two MOUs, NSI
presentations, and a Network Topics list.
UKMO data for UARS Day 2263 and 2264 have been requested but not yet received.
Only remaining UKMO day not received from last weeks outage is 2390.
The CDHF will load February 1998 current day DAAC data in April 1998.
Pending CDHF software upgrades: IDL 5.0, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 20 March 1998 (Orbit 35644, GMT Day 079) through 27
March 1998 (Orbit 35644, GMT Day 086).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 080/2333 35659 Off
HALOE 084/2232 35719 Off
MLS 085/0020 35720 On
HRDI 086/2348 35749 Off
The beta angle is now increasing toward a minimum angle of 0.0 degrees on
02 April 1998 (DOY 092).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.4 degrees C. Additional battery performance data for
each battery on 21 March 1998 and 27 March 1998 are provided in Appendix D.
The clock error ranged from -10.5 to +9.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HALOE 01
% HRDI 01
% MLS 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/041 - 2287.497104
Transponder B: 98/042 - 2287.499330
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/085 show the observatory data loss to be
20 hours, 42 minutes, 33.140 seconds (an increase of 16.480 seconds). This
is a 0.0363 percent data loss which equals a 99.9637 percent data capture
for the mission.
The increase of 16.480 seconds occurred over three days as follows:
A 1.088 second loss on Day 079 due to a hit in the data,
A 10.336 second loss on Day 080 due to insufficient overlap, and
A 4.256 second loss on Day 081 due to hit in the data and on the
transmission line.
Building 23 power was restored as scheduled on Monday, 23 March 1998. No
problems were observed. Operation during the GSFC Building 23 weekend power
outage was nominal with all alternate operational methods successful.
Replay of the telemetry data recorded at the TDRSS White Sands Ground
Terminal was completed and level-0 files created by Friday, 27 March.
No IP testing was executed by the FOT this week. The FOT will resume IP
testing next week, particularly concentrating on Q-channel OBC dumps, and
possible forward link testing. A decision has still not been made by the
UARS FOT and project whether or not to begin testing the IP forward link
directly with the spacecraft (using NOOP commands), or to try to configure
some kind of special test scenario using the UTTS. A deadline for complete
switch-over to the IP network has been set for mid July 1998.
The FOT began making modifications to the OBCSTF resident POCC database.
These changes will be incorporated into the next operational database
release, version 66.
Operational processing of level-0 telemetry analysis files has been
suspended until the UMPGAL cluster and network interface configurations are
finalized.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued.
Continued database work for new PMON software. Worked to define new RTS
and PDB assignments to support PEM day-only operation. These new RTSs and
PDBs will be included with the new PMON software uplink.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 02 April 1998.
A second building 23 power outage is scheduled from about 12 pm, 24 April
1998, until about 12 pm, 27 April 1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS batteries and their performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
Craig Woodruff will no longer be the UARS prime point of
contact in Flight Dynamics Facility (FDF), starting next week. Craig's
responsibilities will now be handled by Dan Andrews.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 March 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
35659 080/2333 98-110 OBC Clk Adj fr 20.6 to 0.00 msec/day
35659 080/2333 98-111 WINDII Power OFF
35685 082/1633 98-112 OBC Clk Adj fr 0.00 to 20.6 msec/DAY
35712 084/1154 98-113 OBC Clk Adj fr 20.6 to 0.00 msec/day
35719 084/2232 98-114 HALOE Power OFF
35720 085/0020 98-115 MLS Power ON
35749 086/2348 98-116 HRDI Low Power mode
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 March 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 35712-19
WIEUTEMP Y-HI WINDII EU Temperature 35721-24
HRPTSPBIMON Y-HI Tele Pulse Bus Cur Monitor 35741
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 March 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
35659 080/23:42 (21 Mar 98) from 20.6 to 00.0 msec/day
35685 082/16:38 (23 Mar 98) from 00.0 to 20.6 msec/day
35712 084/11:58 (25 Mar 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 March 1998
21 Mar 1998 - GMT Day 080 Beta = 43.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +33.6/ -22.4 +6.44/+4.83 +28.8/-16.4 25.0 1.023 22.4
3 +100.8/ +0.0 +5.15/+3.87 +28.0/-16.4 24.8 1.025 21.8
27 Mar 1998 - GMT Day 086 Beta = 22.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 18.4 0.000 0.0
2 +78.4/ -5.6 +7.74/+5.79 +32.0/-16.0 24.0 1.019 24.0
3 +78.4/ +0.0 +6.11/+4.51 +31.2/-16.8 23.8 1.021 23.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 March 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-009 FOT 35631 080/0250 14171 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-012 FOT 35595 076/16:22 19757 Late Acq FOT Conf
98-013 FOT 35603 077/06:06 Tac 4 Crash
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-009 Generic Late Acq for March 1998. Open until 31 March 1998.
SEVNTFW13